The temperature of a body when placed in a room fell from 37⁰C to 35⁰C in 2 minutes. What is the cooling rate?


1
Expert's answer
2022-04-22T13:52:57-0400

The cooling rate

"\frac{\Delta T}{\Delta t}=\rm \frac{(35-37)^{\circ}C}{2\: min}=-1^\circ C/min"